18. On January 1, 2018, Harsel Company issued $400,000, 10-year bonds for $450,000. The bonds pay interest on June 30 and December 31 using a coupon rate is 7%. Using straight line amortization, determine the carrying value of the bonds on December 31, 2018.
A.
$442,500
B.
$447,500
C.
$445,000
D.
$440,000

the answer is C, why?
Answer #1
Premium on issue 50000 =450000-400000
Annual premium amortization 5000 =50000/10
Carrying value of the bonds on December 31, 2018 445000 =450000-5000
Option C $445,000 is correct